Use the elimination method to solve the system.
5a + 8b = 12
11a − 3b = 47

Answer by mananth(16946)

5 a + 8 b = 12 .............1

11 a -3 b = 47 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 8
15 a 24 b = 36
88 a -24 b = 376
Add the two equations
103 a = 412
/ 103
a = 4
plug value of a in (1)
5 a + 8 b = 12
20 + 8 b = 12
8 b = 12 -20
8 b = -8
b = -1 1
x= 4
y= -1
m.ananth@hotmail.ca